OpenStudy (anonymous):

Does anyone know how to graph y=x+1? I don't know how to do this because I am just learning this.

OpenStudy (kc_kennylau):

You may want to find some (x,y) pairs that fit in the equation

OpenStudy (kc_kennylau):

Find y when: (a) x=-1 (b) x=0 (c) x=2

OpenStudy (anonymous):

I have a question that tells me just to graph this and tell weather it is linear or non linear

OpenStudy (kc_kennylau):

I'm creating sub-questions to help you.

OpenStudy (kc_kennylau):

To judge whether it's linear, look at the degree.

OpenStudy (kc_kennylau):

There's no x^2 term or x^3 or any complicated stuff like that

OpenStudy (kc_kennylau):

There's just x and y, no other stuff

OpenStudy (kc_kennylau):

So it's linear

OpenStudy (anonymous):

I am sorry I am not understanding what you are telling me.

OpenStudy (anonymous):

Plug in any value for x. let say x=0 y=x+1 y=0+1 y=1 So we know that when x=0, y=1. hence we have 1 coordinate that lies on the line, which is (0,1). Let's do this again, .. choose a different number lets say x=2 y=x+1 y=2+1 y=3 So we know that when x=2, y=3. hence we have another coordinate that lies on the line, which is (2,3). graph 2 points, (0,1) and (2,3) and connect them with the line. 

OpenStudy (kc_kennylau):

It'd be better to use three points because you don't know if it's linear. Then you might say if it's not linear you can't judge with three points, but well yeah.

OpenStudy (anonymous):

@kc_kennylau she is just starting to graph those... she would get to that later.

OpenStudy (kc_kennylau):

okay...

OpenStudy (anonymous):

I am confused sorry

OpenStudy (kc_kennylau):

Find y when: (a) x=-1 (b) x=0 (c) x=2

OpenStudy (kc_kennylau):

Do you know how to name a point using x- and y-coordinates?

OpenStudy (anonymous):

@fluteprincess352 look at what I wrote.. 1) Do you see how I plugged 0 instead of x, and got the coordinate (0,1) ? 2) Do you see how I plugged 2 instead of x, and got the coordinate (2,3) ? 3) Do you know how to put a coordinate like (1,3) (3,4) and such.. on the graph ? 4) Do you see how I connected the 2 points with the line?

OpenStudy (anonymous):

1. yes 2. yes 3. I think so 4. yes

OpenStudy (kc_kennylau):

Then why are you confused? :O

OpenStudy (anonymous):

I don't know lol! I am terrible at math.

OpenStudy (anonymous):

To verify number 3, can you graph (2, -2) and (-1,3) on the coordinate plane? draw the 2 and say which one is which.  (you can copy and use this picture too)